OpenStudy (anonymous):

since pure gold is soft, gold jewellery is usually made with an alloy. An 18 karat gold alloy contains jewellery 75%(m.m) gold. How much of this alloy can a jeweller make with 8.00g of pure gold can someone help. This is what I did. I know that 18karates is equal to 3.6 g because 18*0.2g=3.6. so gsolvent=75%*3.6g/100% =2.70g and then I added that to the 8g, and get 10.7, but the answer at the back of my text book is 11 g. Did I do something wrong?

OpenStudy (anonymous):

May be its 10.7 Approx 11 . Just round off ..

OpenStudy (anonymous):

okay, I just don't know if I did the question right. :/

OpenStudy (anonymous):

well your calculation looks right to me .

OpenStudy (anonymous):

But I'm not 100 % sure .Hope someone ales help you soon ^^

OpenStudy (aaronq):

Your answer is right, but i think you arrived at it erroneously. I don't understand why you would do this: "18*0.2g=3.6". you can solve this using a ratio though, since 8.00g = 75%, x=100% \(\dfrac{8.00~g}{0.75}=\dfrac{x~g}{1}\rightarrow x=10.666~g\)

OpenStudy (anonymous):

Oh okay thank you. :)